OpenStudy (anonymous):

he cost of running a small factory is modeled by the function C(x) = x^3 - 3 x^2 + 17 x i found that r(x) = 41x but how do i found the p profit =P(x)?

OpenStudy (anonymous):

the cost of running a small factory is modeled by the function C(x) = x^3 - 3 x^2 + 17 x i found that r(x) = 41x but how do i find the profit =P(x)?

OpenStudy (anonymous):

P(x)=R(x)-C(x)

OpenStudy (anonymous):

r(x) is revenue, correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so 41x-( x^3-3x^2+17x)?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

i got -x^3+3x+24

OpenStudy (anonymous):

p(x)=-x^3+3x+24

OpenStudy (anonymous):

p(x)=41x-(x^3-3x^2+17x) p(x)=-x^3+3x^2+24x

OpenStudy (anonymous):

oh yeah the x

OpenStudy (anonymous):

how do i maximize the profit? from that equation

OpenStudy (anonymous):

take the derivative, find the max

OpenStudy (anonymous):

deriv of the profit equation?

OpenStudy (anonymous):

the 2nd deriv test to solve the max parabo;a

OpenStudy (anonymous):

their asking me to Find the number of routers sold which will maximize weekly profit.

OpenStudy (anonymous):

so what's the derivative of p(x)?

OpenStudy (anonymous):

-3x^2+6x+24 or 3x^2+6x+23

OpenStudy (anonymous):

the 1st one

OpenStudy (anonymous):

ok the negative 1 confused me

OpenStudy (anonymous):

so set it to 0

OpenStudy (anonymous):

for the first derivative test, yes

OpenStudy (anonymous):

im tryin to factor out -3(x+3)(x-8)=0

OpenStudy (anonymous):

is it right so far?

OpenStudy (anonymous):

i dont know how to find the vertex (the algebra way)

OpenStudy (anonymous):

-3(x^2-2x-8)

OpenStudy (anonymous):

-3(x-4)(x+2)

OpenStudy (anonymous):

so x= 4 , -2

OpenStudy (anonymous):

yes, can't sell negative number of routers, so -2 is garbage

OpenStudy (anonymous):

2nd derivative is?

OpenStudy (anonymous):

and plug that back into the original equation

OpenStudy (anonymous):

no wait 2nd deriv

OpenStudy (anonymous):

30?

OpenStudy (anonymous):

hmm why the hell is this not working

OpenStudy (anonymous):

hmm why the hell is this not working

OpenStudy (anonymous):

p(x)=-x^3+3x^2+24x p'(x)=-3x^2+6x+24 -3(x^2-2x-8)=0 x^2-2x-8=0 (x-4)(x+2)=0 x=4,-2 Oh, okay, i kept using 3x instead of 3x^2

OpenStudy (anonymous):

p''(x)=-6x+6 p''(4)=-6(4)+6=-24+6=-18 f''(4)<0, maximum at x=4

OpenStudy (anonymous):

plug 4 into original equation